Find the slope and the y-intercept
5 - 1/4 y= 4x
The slope is ____
The y-intercept is (0,__)

Answer by nerdybill(5404)

Find the slope and the y-intercept
5 - 1/4 y= 4x
.
The idea is to "manipulate" the given equation so that it is in the "slope-intercept" form of a line:
y = mx + b
where
m is the slope
b is the y-intercept at (0, b)
.
Starting with:
5 - 1/4 y= 4x
multiplying both sides by 4:
20 - y= 16x
subtracting 20 from both sides:
-y= 16x - 20
multiplying both sides by -1:
y = -16x + 20
.
Therefore,
m = -16 (slope is -16)
b = 20 (y-intercept is at (0,20))